Circular No. 7147

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 H2 (SOHO) D. A. Biesecker, SM&A Corporation and Goddard Space Flight Cen-

ter, reports observations of another probable Kreutz sungrazing comet in SOHO/LASCO C2 data. The comet reached a peak apparent magnitude of ∼ 6.0, and a short tail is evident. It was not visible in the C3 coronagraph because of the vignetting caused by the pylon that holds the solar occulter.

1999 UT  α_{2000}  δ_{2000}
Apr. 19.229  1^{h}53^{m}5  +10^{o}03′

Astrometric measurements, made by Biesecker and reduced by B. G. Marsden, and elements by Marsden are given on MPEC 1999-H07.

COMET C/1999 H1 (LEE)

Orbital elements from MPEC 1999-H06:

T = 1999 July 11.421 TT  ω = 39.^{o}678
Ω = 161.967 }2000.0
q = 0.71496 AU  i = 149.515
1999TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Apr. 16 13^{h}58.^{m}98 −68^{o}23.^{′}6 0.967 1.718 121.^{o}4 29.^{o}9 9.3
18 13 26.07 −68 06.5 0.924 1.690 122.4 30.1 9.1
20 12 51.73 −67 17.7 0.884 1.662 123.2 30.4 8.9
22 12 17.71 −65 52.8 0.847 1.633 123.5 30.9 8.8
24 11 45.64 −63 49.9 0.814 1.604 123.4 31.6 8.6
26 11 16.67 −61 09.9 0.786 1.576 122.7 32.5 8.5
28 10 51.33 −57 56.1 0.762 1.547 121.4 33.7 8.3
30 10 29.62 −54 13.2 0.744 1.518 119.5 35.3 8.2
May 2 10 11.23 −50 07.3 0.732 1.490 117.0 37.1 8.1
4 9 55.73 −45 45.0 0.725 1.461 113.9 39.1 7.9
6 9 42.69 −41 13.2 0.724 1.432 110.4 41.3 7.9
8 9 31.68 −36 38.3 0.729 1.403 106.6 43.6 7.8
10 9 22.35 −32 06.2 0.739 1.375 102.6 45.8 7.7
12 9 14.41 −27 41.9 0.754 1.346  98.4 47.9 7.7

In addition to the discovery information on IAUC 7144, S. Lee notes that the discovery telescope was a 0.41-m f/6 Newtonian reflector (∼ 75×). Visual m_{1} estimates by D. A. J. Seargent, The Entrance, N.S.W.: Apr. 18.45 UT, 9.0 (25×100 binoculars); 19.45, 8.4 (10×50 binoculars).

1999 April 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green